Supplemental Operations Report 13May2023

[title Supplemental Operations Report – May 13th]

Supplemental Operations Report 13May2023

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and installation is pending.

Power system:
Solar: nominal. Input breaker of one of MPTT was off today. After resetting the breaker and solar charger controller current work is nominal.SOC Last 24 hours: Max 100%; Min 41 %; Avg 70.2%.
VDC Last 24 hours: Max 60.36V; Min 43.94V; Avg 50.34V.Generator run time: average run time 3.2 h/day . Generator hours 3810.3. Maintained on May 8th.

Propane Reading Station Tank: 75%
Propane Reading Director Tank: 60%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 72%

Water (Static Tank) – 540 gallons
Water in GreenHab – 190 gallons
Water (Outpost tank) – 380 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: No.
Hours: 191.7
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. To Hanksville for water runs and supplies pick up.

Crew Car used and why, where? No. In Grand Junction.

General notes and comments: Nominal.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Nominal

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Kitchen outlets are not working, inspection will be done on SUnday morning. Septic tank emptied on May 4th, per vendor it was overfield.

Summary of GreenHab Operations: Nominal, heater turned off.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues:

ew parts for Robotic observatory are received and installation is pending.

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Septic tank emptied on May 4th.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site